# A Near-optimal, Scalable and Corruption-tolerant Framework for Stochastic Bandits: From Single-Agent to Multi-Agent and Beyond

### v1

@@ Abstract
We investigate various stochastic bandit problems in the presence of adversarial corruption. A seminal contribution to this area is the BARBAR (Gupta et al., 2019) algorithm, which is both simple and efficient, tolerating significant levels of corruption with nearly no degradation in performance. However, its regret upper bound exhibits a complexity of $O(KC)$, while the lower bound is $\Omega(C)$. In this paper, we enhance the BARBAR algorithm by proposing a novel framework called BARBAT, which eliminates the factor of $K$ and achieves an optimal regret bound up to a logarithmic factor. We also demonstrate how BARBAT can be extended to various settings, including graph bandits, combinatorial semi-bandits, batched bandits and multi-agent bandits. In comparison to the Follow-The-Regularized-Leader (FTRL) family of methods, which provide a best-of-both-worlds guarantee, our approach is more efficient and parallelizable. Notably, FTRL-based methods face challenges in scaling to batched and multi-agent settings.
@@ 1 Introduction
The multi-armed bandit (MAB) problem is one of the most fundamental online learning tasks with partial information feedback, and it has a long and rich history (Lai and Robbins, 1985; Auer et al., 2002). Recently, as security concerns have grown, many studies focus on the so called adversarial corruption regime where an adversary is allowed to manipulate reward observations to an otherwise stochastic environment. Existing approaches can be broadly categorized into two classes: the Follow-The-Regularized-Leader (FTRL) family of algorithms (Zimmert and Seldin, 2021; Ito, 2021b; Jin et al., 2024; Tsuchiya et al., 2024) and elimination-based algorithms (Lykouris et al., 2018; Gupta et al., 2019; Liu et al., 2021; Lu et al., 2021). FTRL-based methods perform optimally in both stochastic and adversarial settings and also achieve optimal regrets in the corruption setting. However, these methods necessitate solving an optimization problem in each round, which may incur high computational costs. Additionally, FTRL-based methods are challenging to extend to parallelizable settings, such as batched bandits and multi-agent bandits. On the other hand, elimination-based methods is simple and computationally efficient. However, whether these methods can achieve optimal regret is still unknown.
@@
In this paper, we improve upon the BARBAR (Gupta et al., 2019) method and propose a novel elimination-based framework called BARBAT (Bad Arms get Recourse, Best Arm gets Trust) for stochastic bandits method and robust to adversarial corruptions. Unlike BARBAR, which employs dynamic epoch lengths, BARBAT adopts static epoch lengths by increasing the probability of selecting the estimated best arm. In addition, BARBAT utilizes epoch-varying failure probabilities $\delta_{m}$ instead of a global failure probability $\delta$ used in BARBAR, allowing us to achieve near-optimal regret bounds. We further demonstrate our BARBAT framework is scalable and parallelizable by extending it to various scenarios, such as the strongly observable graph bandits (Ito et al., 2022; Dann et al., 2023), $d$-set semi-bandits (Zimmert et al., 2019; Ito, 2021a; Tsuchiya et al., 2023), batched bandits (Perchet et al., 2016; Gao et al., 2019; Esfandiari et al., 2021) and cooperative multi-agent multi-armed bandits (CMA2B) (Wang et al., 2022; Liu et al., 2021; Ghaffari et al., 2024). The regret bounds of our methods along with comparisons to recent works are summarized in Table 1.
We summarized our contribution as follows:
@@
  • BARBAT framework. We propose the BARBAT framework for stochastic bandits with adversarial corruptions. Notably, BARBAT resolves the open problem raised by Gupta et al. (2019): eliminating the multiplicative factor $K$ from the corruption-dependent term and achieves near-optimal regret bounds. Moreover, BARBAT does not require prior knowledge of time horizon $T$.
@@
  • Extensions to various corrupted bandit settings. We extend our BARBAT framework to strongly observable graph bandits, $d$-set semi-bandits and batched bandits, demonstrating that all resulting algorithms achieve near-optimal regret. To the best of our knowledge, our BB-BARBAT algorithm for corrupted batched bandits is the first robust algorithm for stochastic batched bandits. We further adapt BARBAT to cooperative multi-agent multi-armed bandits (CMA2B) with adversarial corruption and propose the MA-BARBAT algorithm, which achieves an optimal $V$-fold reduction in regret of each agent. Our MA-BARBAT algorithm only requires $O(V\log(T))$ communication cost.
@@
  • Comparison with FTRL. Compared to FTRL-based methods, our framework has advantages in three folds. First, our methods are more efficient. To compute sample probabilities, FTRL-based methods require iterative processes in each round to find an approximate solution to a constrained convex optimization problem, which can be costly, particularly in semi-bandit settings. Second, our BARBAT framework can be extended to batched and multi-agent settings, whereas extending FTRL to these contexts is quite challenging. Lastly, as shown in the last column of Table 1, our methods do not require the assumption of a unique optimal action, which is required by FTRL algorithms in environments other than MAB.
Technique
Notice that in the BARBAR (Gupta et al., 2019) algorithm, the arm $k$ is approximately drawn $n_{k}^{m}=O(1/\tilde{\Delta}_{k}^{2})$ times in expectation during epoch $m$, where $\tilde{\Delta}_{k}$ is the estimated gap computed in the previous epoch. Thus each epoch having a length of approximately $N_{m}\approx O(\sum_{k}1/\tilde{\Delta}_{k}^{2})$. Consequently, the adversary can significantly extend the length of an epoch by utilizing all corruption budget in the previous epoch, thereby implicitly increasing the number of pulls for suboptimal arms. On the other hand, our BARBAT sets the epoch length of epoch $m$ to be $N_{m}\approx O(K2^{2(m-1)})$ which cannot be affected by the adversary. We still pull each arms approximately $n_{k}^{m}=O(1/\tilde{\Delta}_{k}^{2})$ times in epoch $m$, with the exception of the arm $k_{m}$ that has the best mean rewards in the previous epoch $m-1$. For the estimated optimal arm $k_{m}$, we allocate all remaining pulls, resulting in approximately $N_{m}-\sum_{k\neq k_{m}}n_{k}^{m}$ pulls dedicated to arm $k_{m}$.
However, pulling the estimated optimal arm $k_{m}$ too much times may lead to additional regret, as arm $k_{m}$ could be suboptimal. To mitigate this additional regret, we adopt a epoch-varying failure probabilities $\delta_{m}\approx O(1/(mK2^{2m}))$ for epoch $m$ rather than using a global failure probability $\delta$. This choice of $\delta_{m}$ also eliminate the need for knowing time horizon $T$. Notice that a concurrent work Ghaffari et al. (2024) also employs static epoch lengths and remove the factor $K$ from the regret bound. However, they adopt a global $\delta$ across all epochs, resulting in a regret of $O(C+K\log^{2}(T)/\Delta)$, which could be much worse than our regret $O\left(C+\sum_{\Delta_{k}>0}\frac{\log^{2}(T)}{\Delta_{k}}\right)$. Here $\Delta_{k}$ is the suboptimality gap of arm $k$ and $\Delta=\min_{\Delta_{k}>0}\Delta_{k}$ is the minimum positive suboptimality gap.
@@
In recent years, the study of stochastic bandits with adversarial corruptions has garnered significant attention. Existing approaches can be broadly classified into two main categories: elimination-based algorithms (Lykouris et al., 2018; Gupta et al., 2019; Liu et al., 2021; Lu et al., 2021) and FTRL-based algorithms (Zimmert and Seldin, 2021; Ito, 2021b; Jin et al., 2024; Tsuchiya et al., 2024). Elimination-based methods, with the Active Arm Elimination (AAE) algorithm (Even-Dar et al., 2006) being a prominent example, progressively discard suboptimal arms based on empirical reward estimates. By removing poorly performing arms, these algorithms focus their sampling efforts on a shrinking set of promising candidates. Lykouris et al. (2018) introduced a multi-layer extension of the AAE algorithm, achieving a regret bound of $\widetilde{O}\left(KC\sum_{\Delta_{k}>0}1/\Delta_{k}\right)$. Building on this, Gupta et al. (2019) proposed the BARBAR algorithm and reduce the regret to $\widetilde{O}\left(KC+\sum_{\Delta_{k}>0}1/\Delta_{k}\right)$. On the other hand, FTRL-based methods (Bubeck and Slivkins, 2012; Seldin and Slivkins, 2014; Auer and Chiang, 2016; Wei and Luo, 2018; Zimmert and Seldin, 2021; Ito, 2021b; Jin et al., 2024; Tsuchiya et al., 2024) can achieve optimal regret bound in both stochastic setting, adversarial setting and the adversarial corruption setting. However, these methods are computationally expensive than elimination-based methods since they need to solve an optimization problem in each round.
Stochastic bandits with adversarial corruptions can be extended to graph bandits, semi-bandits, and batched bandits, each of which benefits from specialized techniques to handle challenges such as exponential action spaces, partial observability, or limited adaptivity (Rouyer et al., 2022; Ito et al., 2022; Dann et al., 2023; Zimmert et al., 2019; Ito, 2021a; Tsuchiya et al., 2023; Perchet et al., 2016; Gao et al., 2019; Esfandiari et al., 2021). Finally, multi-agent extensions (Liu et al., 2021; Wang et al., 2022; Chawla et al., 2023; Ghaffari et al., 2024) enable agents to pool information and expedite the learning process but often encounter larger regret bounds under adversarial corruptions or rely on group-level performance guarantees rather than individual regrets.
@@
Multi-arm bandits Zimmert and Seldin (2021) $C+\sum_{\Delta_{k}>0}\frac{\log(T)}{\Delta_{k}}$
Jin et al. (2024) $C+\sum_{\Delta_{k}>0}\frac{\log(T)}{\Delta_{k}}$
@@
Strongly observable graph bandits Lu et al. (2021) $\alpha C+\sum_{k\in{\mathcal{D}}^{*}}\frac{\log^{2}(T)}{\Delta_{k}}$
Ito et al. (2024) $C+\frac{\alpha(\ln(K/\alpha+1))\log(T)}{\Delta}$
Ito et al. (2022) $C+\frac{\alpha\log^{3}(T)}{\Delta}$
Dann et al. (2023) $C+\frac{\min\{\widetilde{\alpha},\alpha\log(K)\}\log(T)}{\Delta}$
@@
$d$-set semi-bandits Zimmert et al. (2019); Ito (2021a) Tsuchiya et al. (2023) $dC+\sum_{\Delta_{k}>0}\frac{\log(T)}{\Delta_{k}}$
@@
Batched bandits Algorithm 5 $CT^{\frac{1}{M+3}}+T^{\frac{4}{M+3}}\Bigl{(}$ $\sum_{\Delta_{k}>0}\frac{M\log(T)}{\Delta_{k}}+\frac{K\log(T)}{M\Delta}\Bigr{)}$
Multi-agent multi-arm bandits Liu et al. (2021) $VC+\frac{K\log^{2}(T)}{\Delta}$
Ghaffari et al. (2024) $\frac{C}{V}+\frac{K\log^{2}(T)}{V\Delta}$
Algorithm 6 $\frac{C}{V}+\sum_{\Delta_{k}>0}\frac{\log^{2}(T)}{V\Delta_{k}}$
Table 1: “Prop. 1” indicates the property that the algorithm can efficiently obtain the sampling probablity of arms. “Prop. 2” indicates the property that the assumption of unique optimal action is not required. $\Delta_{k}$ denote the suboptimality gap of arm $k$, and let $\Delta$ is the smallest positive suboptimality gap. For graph bandits, $\alpha$ denotes the independence number of the feedback graph $G$ and ${\mathcal{I}}^{*}$ denotes the set of at most $O\!\Bigl{(}\alpha\ln\bigl{(}\frac{K}{\alpha}+1\bigr{)}\Bigr{)}$ arms with the smallest gap. $\dagger$": these works only consider undirected graphs while rest works for graph bandits consider directed graphs. “*": the definition of $d$-set semi-bandits is introduced in Section 4.2. Here, we assume without loss of generality that $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{K}$. and overloading the notation $\Delta_{k}=\mu_{k}-\mu_{d}$ for each arm $k>d$.
@@ 2 Preliminaries
We consider stochastic multi-armed bandits with adversarial corruptions. In this setting, the agent interacts with the environment over $T$ rounds by selecting an arm from a set of $K$ arms, denoted by $[K]$. In each round $t$, the environment generates a random reward vector $\{r_{t,k}\}_{k\in[K]}$. An adversary, having access to the reward vector, subsequently attack these rewards to produce the corrupted reward vector $\{\widetilde{r}_{t,k}\}_{k\in[K]}$. The agent then selects an arm $I_{t}$ according to its strategy and observes the corresponding corrupted reward $\widetilde{r}_{t,I_{t}}$. Let $\mu_{k}$ denote the mean reward of arm $k\in[K]$, and let $k^{*}\in\arg\max_{k\in[K]}\mu_{k}$ be an optimal arm. The suboptimality gap for arm $k$ is defined as $\Delta_{k}=\mu_{k}-\mu_{k^{*}}$, and we denote $\Delta=\min_{\Delta_{k}>0}\Delta_{k}$ as the smallest suboptimality positive gap. The corruption level is define as $C=\sum_{t=1}^{T}\max_{k\in[K]}\left|\widetilde{r}_{t,k}-r_{t,k}\right|$. Our goal is to minimize the expectation of the pseudo-regret $R(T)$
$R(T)=\sum_{t=1}^{T}\mu_{k^{*}}-\sum_{t=1}^{T}\mu_{I_{t}}=\sum_{t=1}^{T}\Delta_{I_{t}}.$
Notice that FTRL-based methods (Rouyer et al., 2022; Ito et al., 2022; Dann et al., 2023; Zimmert et al., 2019; Ito, 2021a; Tsuchiya et al., 2023; Perchet et al., 2016; Gao et al., 2019; Esfandiari et al., 2021) usually consider another pseudo-regret $\widetilde{R}(T)=\max_{k}{\mathbb{E}}\Bigl{[}\sum_{t=1}^{T}\bigl{(}r_{t,I_{t}}-r_{t,k}\bigr{)}\Bigr{]}$. Theorem 3 in Liu et al. (2021) presents the conversion between the two definition of pseudo-regrets in the adversarial corruption as ${\mathbb{E}}\bigl{[}R(T)\bigr{]}=\Theta\bigl{(}\widetilde{R}(T)+O(C)\bigr{)}$.

### v2

# A Near-optimal, Scalable and Parallelizable Framework for Stochastic Bandits Robust to Adversarial Corruptions and Beyond

@@ Abstract
We investigate various stochastic bandit problems in the presence of adversarial corruptions. A seminal work for this problem is the BARBAR gupta2019better algorithm, which achieves both robustness and efficiency. However, it suffers from a regret of $O(KC)$, which does not match the lower bound of $\Omega(C)$, where $K$ denotes the number of arms and $C$ denotes the corruption level. In this paper, we first improve the BARBAR algorithm by proposing a novel framework called BARBAT, which eliminates the factor of $K$ to achieve an optimal regret bound up to a logarithmic factor. We also extend BARBAT to various settings, including multi-agent bandits, graph bandits, combinatorial semi-bandits and batched bandits. Compared with the Follow-the-Regularized-Leader framework, our methods are more amenable to parallelization, making them suitable for multi-agent and batched bandit settings, and they incur lower computational costs, particularly in semi-bandit problems. Numerical experiments verify the efficiency of the proposed methods.
@@ 1 Introduction
The multi-armed bandit (MAB) problem is one of the most fundamental online learning tasks, and it has a long and rich history lai1985asymptotically; auer2002finite. Recently, as security concerns have grown, many studies focus on the so-called adversarially corrupted regime where an adversary is allowed to manipulate reward observations in a stochastic environment. Existing approaches can be broadly categorized into two classes: the Follow-the-Regularized-Leader (FTRL) family of algorithms zimmert2021tsallis; ito2021parameter; jin2023improved; tsuchiya2023stability and elimination-based algorithms lykouris2018stochastic; gupta2019better; liu2021cooperative; lu2021stochastic. FTRL-based methods perform optimally in both stochastic and adversarial settings and also achieve optimal regret in the adversarially corrupted setting. However, these methods need to solve an optimization problem in each round, which may incur high computational costs in many scenarios such as semi-bandits. Also, the FTRL framework is challenging to extend to parallel settings such as batched bandits and multi-agent bandits. On the other hand, elimination-based methods such as BARBAR gupta2019better are computationally efficient and easy to parallelize. However, the BARBAR algorithm suffers from a suboptimal regret of $O(KC)$, and it remains an open question how to achieve optimal regret by elimination-based methods.
@@
In this paper, we improve upon the BARBAR gupta2019better method and propose a novel framework called BARBAT (Bad Arms get Recourse, Best Arm gets Trust) for stochastic bandits with adversarial corruptions. Unlike BARBAR, which employs dynamic epoch lengths, BARBAT adopts static epoch lengths by increasing the probability of selecting the estimated best arm. In addition, BARBAT utilizes epoch-varying failure probabilities $\delta_{m}$ instead of a global failure probability $\delta$ used in BARBAR, allowing us to achieve near-optimal regret bounds. We further demonstrate that our BARBAT framework is scalable and parallelizable by extending it to various scenarios, including cooperative multi-agent multi-armed bandits (CMA2B) wang2023achieving; liu2021cooperative; ghaffari2025multi, graph bandits ito2022nearly; dann2023blackbox, semi-bandits zimmert2019beating; ito2021hybrid; tsuchiya2023further and batched bandits perchet2015batched; gao2019batched; esfandiari2021regret. The regret bounds of our methods along with comparisons to recent works are summarized in Table 1.
We summarize the contributions of this work as follows:
@@
  • BARBAT framework  We propose the BARBAT framework for stochastic bandits with adversarial corruptions that achieves a near-optimal regret bound. This result resolves the open problem raised in gupta2019better. Moreover, BARBAT does not require prior knowledge of the time horizon $T$.
@@
  • Extensions of BARBAT  We extend our BARBAT framework to various corrupted bandit settings including multi-agent bandits, batched bandits, strongly observable graph bandits and $d$-set semi-bandits. We show that all resulting algorithms achieve near-optimal regret. To the best of our knowledge, our work is the first to study a robust algorithm for batched bandits. We also provide a lower bound for batched bandits with adversarial corruptions.
@@
  • Comparison with FTRL  Compared to FTRL-based methods, our framework has advantages in three ways. First, our methods are more efficient. The FTRL framework requires solving a constrained convex optimization problem in each round. Such problems generally do not have a closed-form solution (except for FTRL with Shannon-entropy regularization) and can therefore be computationally costly, particularly in semi-bandit scenarios. Second, our BARBAT framework is parallelizable, making it suitable for batched and multi-agent settings, whereas extending FTRL to these scenarios is quite challenging. Lastly, as shown in the last column of Table 1, our methods do not require the assumption of a unique optimal action, which is required by FTRL-based methods in environments other than MAB.
@@
In recent years, the study of stochastic bandits with adversarial corruptions has garnered significant attention. Existing approaches can be broadly classified into two main categories: elimination-based algorithms (lykouris2018stochastic; gupta2019better; liu2021cooperative; lu2021stochastic) and FTRL-based algorithms (zimmert2021tsallis; ito2021parameter; jin2023improved; tsuchiya2023stability). Elimination-based methods, with the Active Arm Elimination (AAE) algorithm (even2006action) being a prominent example, progressively discard suboptimal arms based on empirical reward estimates. By removing poorly performing arms, these algorithms focus their sampling efforts on a shrinking set of promising candidates. lykouris2018stochastic introduced a multi-layer extension of the AAE algorithm, achieving a regret bound of $\widetilde{O}\left(KC\sum_{\Delta_{k}>0}1/\Delta_{k}\right)$. Building on this, gupta2019better proposed the BARBAR algorithm and reduced the regret to $\widetilde{O}\left(KC+\sum_{\Delta_{k}>0}1/\Delta_{k}\right)$. On the other hand, FTRL-based methods (bubeck2012best; seldin2014one; auer2016algorithm; wei2018more; zimmert2021tsallis; ito2021parameter; jin2023improved; tsuchiya2023stability) can achieve optimal regret bounds in the stochastic, adversarial and adversarially corrupted settings. However, these methods are more computationally expensive than elimination-based methods since they need to solve an optimization problem in each round.
Stochastic bandits with adversarial corruptions can be extended to graph bandits, semi-bandits, and batched bandits, each of which benefits from specialized techniques to handle challenges such as exponential action spaces, partial observability, or limited adaptivity (rouyer2022near; ito2022nearly; dann2023blackbox; zimmert2019beating; ito2021hybrid; tsuchiya2023further; perchet2015batched; gao2019batched; esfandiari2021regret). Finally, multi-agent extensions (liu2021cooperative; wang2023achieving; chawla2023collaborative; ghaffari2025multi) enable agents to pool information and expedite the learning process but often encounter larger regret bounds under adversarial corruptions or rely on group-level performance guarantees rather than individual regrets.
@@
Multi-arm bandits BARBAR gupta2019better $KC+\sum_{\Delta_{k}>0}\frac{\log^{2}(T)}{\Delta_{k}}$
$\frac{1}{2}$-Tsallis-FTRL zimmert2021tsallis $C+\sum_{\Delta_{k}>0}\frac{\log(T)}{\Delta_{k}}$
$\frac{1}{2}$-Tsallis-FTRL jin2023improved $C+\sum_{\Delta_{k}>0}\frac{\log(T)}{\Delta_{k}}$
Shannon-FTRL jin2023improved; ito2022nearly $C+\sum_{\Delta_{k}>0}\frac{\log^{2}(T)}{\Delta_{k}}$
@@
Lower bound gupta2019better $C+\sum_{\Delta_{k}>0}\frac{\log(T)}{\Delta_{k}}$
Multi-agent multi-arm bandits CBARAC liu2021cooperative $C+\frac{K\log^{2}(T)}{\Delta}$
DRAA ghaffari2025multi $\frac{C}{V}+\frac{K\log^{2}(T)}{V\Delta}$
Algorithm 2 $\frac{C}{V}+\sum_{\Delta_{k}>0}\frac{\log^{2}(T)}{V\Delta_{k}}$
Lower bound $\frac{C}{V}+\sum_{\Delta_{k}>0}\frac{\log(T)}{V\Delta_{k}}$
Batched bandits Algorithm 3 $CT^{\frac{1}{L+3}}+T^{\frac{4}{L+3}}\Big($ $\sum_{\Delta_{k}>0}\frac{L\log(T)}{\Delta_{k}}+\frac{K\log(T)}{L\Delta}\Big)$
Lower bound (Ours) $T^{1/L}(K+C^{1-1/L})$
Strongly observable graph bandits Elise lu2021stochastic $\alpha C+\sum_{k\in{\mathcal{D}}^{*}}\frac{\log^{2}(T)}{\Delta_{k}}$
Shannon-FTRL ito2022nearly $C+\frac{\alpha\log^{3}(T)}{\Delta}$
$(1{-}\frac{1}{\log(K)})$-Tsallis-FTRLdann2023blackbox $C+\frac{\min\{\widetilde{\alpha},\alpha\log(K)\}\log(T)}{\Delta}$
@@
Lower bound $C+\sum_{k\in{\mathcal{I}}^{*}}\frac{\log(T)}{\Delta_{k}}$
$d$-set semi-bandits HYBRID zimmert2019beating $dC+\sum_{\Delta_{k}>0}\frac{\log(T)}{\Delta_{k}}$
LBINF (ito2021hybrid) $dC+\sum_{\Delta_{k}>0}\frac{\log(T)}{\Delta_{k}}$
LBINF_LS, LBINF_GD tsuchiya2023further $dC+\sum_{\Delta_{k}>0}\frac{\log(T)}{\Delta_{k}}$
@@
Lower bound $dC+\sum_{\Delta_{k}>0}\frac{\log(T)}{\Delta_{k}}$
Table 1: “Prop. 1" denotes that the algorithm can efficiently compute the sampling probabilities of the arms, while “Prop. 2” denotes that the assumption of a unique optimal arm is not required. For graph bandits, $\alpha$ denotes the independence number of graph $G$ and $\mathcal{I}$$^{*}$ denotes the set of at most $O\big(\alpha\ln\bigl(\frac{K}{\alpha}+1\bigr)\big)$ arms with the smallest gaps. For batched bandits, $L$ denotes the number of batches.
$\dagger$": Elise is restricted to undirected graphs, while the remaining studies address directed graphs.
$\ddagger$": These lower bounds can be directly achieved by combining the corruption lower bound from gupta2019better with the lower bound for the stochastic setting.
@@ 2 Preliminaries
We consider stochastic multi-armed bandits with adversarial corruptions. In this setting, the agent interacts with the environment over $T$ rounds by selecting an arm from a set of $K$ arms, denoted by $[K]$. In each round $t$, the environment generates an i.i.d. random reward vector $\{r_{t,k}\}_{k\in[K]}$ from an unknown fixed distribution. An adversary, having access to the reward vector, subsequently attacks these rewards to produce the corrupted reward vector $\{\widetilde{r}_{t,k}\}_{k\in[K]}$. The agent then selects an arm $I_{t}$ according to its strategy and observes the corresponding corrupted reward $\widetilde{r}_{t,I_{t}}$. Let $\mu_{k}$ denote the mean reward of arm $k\in[K]$, and let $k^{*}\in\arg\max_{k\in[K]}\mu_{k}$ be an optimal arm. The corruption level is defined as $C=\sum_{t=1}^{T}\max_{k\in[K]}\left|\widetilde{r}_{t,k}-r_{t,k}\right|$. The suboptimality gap for arm $k$ is defined as $\Delta_{k}=\mu_{k}-\mu_{k^{*}}$, and we denote $\Delta=\min_{\Delta_{k}>0}\Delta_{k}$ as the smallest positive suboptimality gap.
Our goal is to minimize the pseudo-regret:
$R(T)=\max_{k\in[K]}{\mathbb{E}}\left[\sum_{t=1}^{T}r_{t,k}\right]-{\mathbb{E}}\left[\sum_{t=1}^{T}r_{t,I_{t}}\right]=\sum_{t=1}^{T}\mu_{k^{*}}-\sum_{t=1}^{T}\mu_{I_{t}}=\sum_{t=1}^{T}\Delta_{I_{t}},$
which is a standard definition in stochastic bandits. Notice that FTRL-based methods (rouyer2022near; ito2022nearly; dann2023blackbox; zimmert2019beating; ito2021hybrid; tsuchiya2023further; perchet2015batched; gao2019batched; esfandiari2021regret) typically adopt a different form of pseudo-regret, denoted as $\widetilde{R}(T)=\max_{k}{\mathbb{E}}\big[\sum_{t=1}^{T}(\tilde{r}_{t,I_{t}}-\tilde{r}_{t,k})\big]$, which is more appropriate for adversarial settings. However, we argue that comparing against the true means, rather than the corrupted rewards, is more natural. Theorem 3 in liu2021cooperative presents the conversion between the two definitions of pseudo-regret in the adversarially corrupted setting as $R(T)=\Theta(\widetilde{R}(T)+O(C))$.
CMA2B
In cooperative multi-agent multi-armed bandits, $V$ agents collaborate to play the bandit game. We denote the set of agents by $[V]$. All agents are allowed to pull any arm from $[K]$. When multiple agents pull the same arm, each receives an independent reward drawn from that arm’s distribution. The adversary is allowed to attack all agents with a total corruption budget $C$. We only consider the centralized setting where each agent can broadcast messages to all other agents without any delays. For simplicity, we define the total communication cost for each agent $v$ as:
[…]
We aim to minimize the individual pseudo-regret for each agent $v$ over a horizon of $T$ rounds:
$R_{v}(T)=\max_{k\in[K]}{\mathbb{E}}\left[\sum_{t=1}^{T}r_{v,t}(k)\right]-{\mathbb{E}}\left[\sum_{t=1}^{T}r_{v,t}({I_{v,t}})\right]=\sum_{t=1}^{T}(\mu_{k^{*}}-\mu_{I_{v,t}})=\sum_{t=1}^{T}\Delta_{I_{v,t}}.$
Batched bandits
In this case, we suppose the agent can only observe the corrupted rewards from a batch after it has concluded. The time horizon $T$ is divided into $L$ batches, represented by a grid ${\mathcal{T}}=\{t_{1},t_{2},\dots,t_{L}\}$, where $1\leq t_{1}<t_{2}<\cdots<t_{L}=T$. As in most previous works gao2019batched, the number of batches $L$ of interest is at most $\log(T)$. Additionally, the grid can be one of two types:
  • Static Grid: The grid is fixed and predetermined before any arm sampling occurs.
  • Adaptive Grid: The value of $t_{j}$ is determined dynamically after observing the rewards up to time $t_{j-1}$, which may incorporate external randomness.
Previous works (perchet2015batched; gao2019batched; zhang2020inference; esfandiari2021regret; jin2021double) mainly focus on stochastic batched bandits. To the best of our knowledge, no existing work studies stochastic batched bandits under adversarial corruptions.
Strongly observable graph bandits
In this setting, when the agent pulls arm $k$, it may observe the rewards of other arms. This reward-feedback structure is represented by a directed graph $G=([K],E)$, where pulling arm $k_{i}$ reveals the rewards of each arm $k_{j}$ such that $(k_{i},k_{j})\in E$. Each arm either has a self-loop or receives incoming edges from all other arms (alon2017nonstochastic). Let $\alpha$ denote the independence number of $G$, i.e., the maximum size of an independent set. A vertex set $D$ is called an out-domination set of $G$ if every vertex in $G$ has an incoming edge from some vertex in $D$.
$d$-set semi-bandits
The $d$-set semi-bandits problem is a special case of semi-bandits. Given $d\in[K{-}1]$, in each round $t$ the agent selects a combinatorial action $\mathbf{x}_{t}$ (containing $d$ distinct arms) from the set $\mathcal{X}=\big\{\mathbf{x}\in\{0,1\}^{K}\,\big|\,\sum_{k=1}^{K}\mathbf{x}(k)=d\big\}$, where $\mathbf{x}(k)=1$ indicates that arm $k$ is selected. Then the agent can observe the corrupted rewards for each arm in the action (i.e., for each arm $k$ such that $\mathbf{x}_{t}(k)=1$). Let $\mu_{k}$ denote the mean reward of arm $k\in[K]$. Without loss of generality, we assume $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{K}$ and denote $\Delta_{k}=\mu_{d}-\mu_{k}$ for all $k>d$. We aim to minimize the pseudo-regret over a horizon of $T$ rounds:
$R(T)=\max_{\mathbf{x}\in{\mathcal{X}}}{\mathbb{E}}\left[\sum_{t=1}^{T}\sum_{k:\mathbf{x}[k]=1}r_{t,k}\right]-{\mathbb{E}}\left[\sum_{t=1}^{T}\sum_{k:\mathbf{x}_{t}[k]=1}r_{t,k}\right]=\sum_{t=1}^{T}\bigg(\sum_{k=1}^{d}\mu_{k}-\sum_{k:\mathbf{x}_{t}(k)=1}\mu_{k}\bigg).$